# Need for zkSpeed: Accelerating HyperPlonk for Zero-Knowledge Proofs

### v1

@@ Abstract.
(Preprint) Zero-Knowledge Proofs (ZKPs) are rapidly gaining importance in privacy-preserving and verifiable computing. ZKPs enable a proving party to prove the truth of a statement to a verifying party without revealing anything else. ZKPs have applications in blockchain technologies, verifiable machine learning, and electronic voting, but have yet to see widespread adoption due to the computational complexity of the proving process. Recent works have accelerated the key primitives of state-of-the-art ZKP protocols on GPU and ASIC. However, the protocols accelerated thus far face one of two challenges: they either require a trusted setup for each application, or they generate larger proof sizes with higher verification costs, limiting their applicability in scenarios with numerous verifiers or strict verification time constraints. This work presents an accelerator, zkSpeed, for HyperPlonk, a state-of-the-art ZKP protocol that supports both one-time, universal setup and small proof sizes for typical ZKP applications in publicly verifiable, consensus-based systems. We accelerate the entire protocol, including two major primitives: SumCheck and Multi-scalar Multiplications (MSMs). We develop a full-chip architecture using 366.46 mm2 and 2 TB/s of bandwidth to accelerate the entire proof generation process, achieving geometric mean speedups of 801 over CPU baselines.
@@ 1. Introduction
Zero knowledge proofs enable a prover to produce a certificate or proof that some computation (possibly with secret inputs) was performed correctly. The proof itself reveals no information about the secret inputs and can be verified much faster than the original computation. Proof systems can be characterized by several important criteria: The prover time, the verifier time, the proof size, and the cryptographic assumptions required for the proof to be secure. Each application of proof systems results in a different set of requirements on these criteria. For instance, private transactions (Ben-Sasson et al., 2014) require a proof per transaction to be posted on a blockchain and distributed to all blockchain nodes, thus prioritizing proof size.
Recent years have seen an explosion in proof systems (Groth, 2016; Gabizon et al., 2019; Chen et al., 2022; Xie et al., 2022). Each of them achieves a different Pareto optimal point with respect to the criteria. In all applications, prover time is important, and fast provers enable new applications. Recently, two proof systems have been accelerated in hardware: Groth16 (Groth, 2016) in pipeZK (Zhang et al., 2021) and SZKP (Daftardar et al., 2024), and Orion with NoCap (Samardzic et al., 2024).
@@
Groth16 produces very short proofs (188 bytes) and enables millisecond verification, independent of computation size. However, it relies on a strong cryptographic assumption; a circuit-specific trusted setup where a trusted party generates keys using secret randomness. If this randomness is not properly discarded or the party is malicious, the system’s security is compromised, allowing false statements to be proven. Hence, Zcash and other applications have moved away from circuit-specific trusted-setup protocols (Nelson, 2022). On the other hand, Orion does not require a trusted setup, has a fast prover, and does not rely on elliptic curve cryptography, but has very large (8MB) proofs. This is 4 larger than the maximum block size of about 2MB in Ethereum cryptocurrency. Thus, a private transaction with an Orion proof would not fit into a single block.
In this work, we accelerate HyperPlonk (Chen et al., 2022), a recent ZKP system gaining attention for its intriguing set of tradeoffs, making it suitable for many applications. Its proofs are about 5 KB and it uses a so-called universal trusted setup (Groth et al., 2018). This is run only once and reused for any computation. It is easy to run the setup among many parties. The security is guaranteed as long as at least one party is honest. Unlike Groth16 and others, it does not use the Number Theoretic Transform (NTT), instead using SumCheck. NTTs asymptotically run in time ( represents the size of underlying computation), while SumCheck runs in linear time.
Accelerating HyperPlonk presents its own challenges. First, as noted, the protocol replaces NTT, well studied in the hardware literature, with the SumCheck protocol, which is relatively less understood from a hardware standpoint. Although implemented in NoCap (Samardzic et al., 2024), as we will shortly discuss, HyperPlonk’s SumChecks are significantly more complex, incurring greater bandwidth costs and design complexity. Further, HyperPlonk introduces additional runtime computations like modular inverses that are not needed in Groth16 and Orion, and thus not studied from a hardware standpoint in this prior body of work. The additional design complexity requires careful architectural design, new hardware optimizations, and thorough design space exploration to identify architectures with area-performance tradeoffs that justify hardware acceleration.
In this paper we present zkSpeed, a modular HyperPlonk accelerator that overcomes the above challenges. zkSpeed comprises eight unique accelerator units that are composed into a complete architecture, including shared and local scratchpads to capture on-chip data reuse. Units and SRAMs communicate via a multi-channel shared bus, and we develop a mapping of complete Hyperplonk protocol onto zkSpeed. This paper makes the following contributions:
@@
A high-throughput, fully-pipelined accelerator to handle three flavors of SumCheck for HyperPlonk-based proofs
@@
A novel implementation for modular inversion to compute fraction polynomials not seen in other ZKP protocols
@@
Additional optimizations across all hardware units that save upwards of 50% area per unit and up to 85% in bandwidth.
@@
A full-chip design that achieves 801 gmean speedup over CPU at iso-compute area
@@ 2. Background
State-of-the-art ZKP protocols are zero-knowledge Succinct Non-interactive Argument of Knowledge (zkSNARKs). zkSNARKs have three properties: (i) zero-knowledge, i.e., the proof does not reveal any information about the secret witness ; (ii) succinct, i.e., the proof has a few hundreds of bytes; and (iii) non-interactive, i.e., sends the proof to in one exchange.
zkSNARKs like HyperPlonk use polynomials to encode the correct execution of the target program. Polynomials can encode complex computations and constraints in a compact form, reducing the computational burden on the to a few checks, instead of requiring inspection of individual operations in the program. In HyperPlonk, the two most time consuming kernels are SumCheck and Multiscalar Multiplications (MSM). The SumCheck kernel lets the demonstrate knowledge of a polynomial by verifying that its properties hold over a large set of inputs without revealing the polynomial. MSMs ensure that is bound to that polynomial, preventing manipulation of the polynomial in the proof.
@@ 2.2. The SumCheck Kernel
SumCheck is an interactive protocol between a prover and a verifier . demonstrates to , that it has correctly computed the sum of a polynomial over the boolean hypercube, i.e., over all Boolean (0/1) assignments of its variables (Thaler, 2022).
@@
and returns it to in the form of its coefficients. checks that , and if so, generates a random challenge and sends it to . asks to prove that
which is an instance of SumCheck, except over a ()-variate polynomial. Thus, in Round 2, computes and returns
@@ 2.3. Multilinear Polynomials
is multilinear because the maximum degree of individual variables is one. ZKP protocols use a general representations of multilinear polynomials, shown below for a 2-variable multilinear polynomial:
These representations map cleanly to hardware, since we can store polynomials as lookup tables indexed by the binary values of . In general, a multilinear polynomial with variables can be stored in a lookup table of entries. As we will discuss in Section 3, HyperPlonk uses multilinear polynomials as building blocks for higher-degree polynomials on which we then perform SumCheck and other computations. In the rest of the paper, we will use the term MLE table to refer to these lookup data structures (MLE stands for ”multilinear extensions” (Thaler, 2022); the word “extension” reflects the fact that these polynomials can also be evaluated at non-binary, or extended, values).
@@ 2.4. The MSM Kernel
MSMs are dot products between a vector of scalars and a vector of 2D or 3D points on an elliptical curve, e.g. . MSMs are used in ZKPs to perform commitments. A commitment is a cryptographic primitive that binds a prover to a value without revealing it. In zkSNARK protocols like HyperPlonk (Chen et al., 2022), the scalars are the entries of the MLE tables. Computing a dot-product reduces these polynomials to a single value, i.e., the commitment.
MSMs are a bottleneck in ZKP provers and recent work has focused on accelerating them on ASIC and GPU (Daftardar et al., 2024; Zhang et al., 2021; Ma et al., 2023; Liu et al., 2024a; Ji et al., 2024; Lu et al., 2022; Zhou et al., 2024). The bottleneck is due to the extremely expensive elliptic curve point multiplications that they use. To reduce this cost, MSMs use Pippenger’s algorithm (Pippenger, 1976), which performs point multiplications via several point additions (PADDs). PADDs are still expensive, typically tens of regular multiplications. In the context of ASIC accelerators, the state-of-the-art (SZKP (Daftardar et al., 2024)) presents a framework for building scalable MSM architectures.
@@ 3.1. Plonk-based encodings
This is the first work to accelerate HyperPlonk, a zkSNARK protocol that replaces Number Theoretic Transform (Fast Fourier Transforms over finite fields) with the SumCheck protocol. In ZKPs, the program being proven must be converted into a specific form before generating the proof. While most prior works use Rank-1 Constraint System (R1CS)—a series of sparse matrix-vector encodings—HyperPlonk adopts a Plonk-based structure (Gabizon et al., 2019), which we describe shortly. R1CS and Plonk encodings map all program values into addition, multiplication, and boolean operations, with nonlinear operations (branches) resolved via bit-wise decompositions.
Plonk-based encodings mapping each operation in a program’s execution to an arithmetic logic unit (or ”gate”) that supports only addition, multiplication, and equality checks, as shown in Eq. 1:
Terms are binary “control” signals for left, right, multiply, and output ports. Terms represent the gate’s inputs, and is its output. represents a constant input. An addition operation, for instance, is implemented by setting , , and other control inputs to . This yields: . That is, if and only if the addition is correctly performed.
Each operation in a program is mapped to a Plonk gate. A program with operations has Plonk gates interconnected to form a “circuit.” A -dimensional polynomial represents the entire circuit, with binary assignments to variables representing individual gates. For example, if , the circuit has gates, and represents the first gate.
is constructed using multilinear polynomials representing every term from Eq. 1. For example, is the left control input polynomial, is the first data input polynomial, and so on. Each of these polynomials also takes as an ”index”. For example, in a program that maps to 16 gates, will tell us whether or not the ”multiply” control signal is enabled in gate 2. tells us what the output data signal is at gate 15. All individual polynomials are encoded in this fashion, and can be stored as MLE tables. These MLE tables are populated from the program trace in software before running the HyperPlonk prover. The MLEs are then used by the steps of the HyperPlonk protocol, which we discuss in the following sections. A consequence is that the SumCheck algorithm (Section 2.2) has to be modified to account for Plonk polynomials.
@@ 3.2. SumCheck on Plonk-based Encodings
The key difference between the SumCheck used in HyperPlonk and the example shown in Section 2.2 is in the polynomial structure. While that example shows how SumCheck runs on a single polynomial of arbitrary degree, HyperPlonk runs SumChecks on the Plonk polynomial, which involves the products of multilinear polynomials. The polynomial in Eq. 1, for example, consists of 5 terms, each of a different degree. Figure 1 shows an example of how a SumCheck of a three-polynomial product is computed.
In the example from Section 2.2, we computed the evaluations at 0 and 1, because the result of summing over all variables yielded a univariate polynomial of degree 1. In this example, however, when we sum over all variables, we have a univariate degree 3 polynomial. From elementary algebra, such a polynomial must be evaluated at 4 unique points to fully characterize the polynomial. In the figure example, in the blue region, we are iterating over to compute the evaluations. For each iteration (red region), we must evaluate each polynomial at . Then, in the green region, we compute the product for across all polynomials, and then sum across iterations. This is repeated for . After the summations, each polynomial’s MLE table is updated with a random challenge from the verifier. For example, a table (for round 2) can be constructed from an table (from round 1) using the formula .
These core computational steps are more expensive than the baseline SumCheck. An additional layer of complexity is that in the polynomial shown in Eq. 1, there are 5 terms of varying degrees. The varying degrees leads to imbalance in how many evaluations are needed (i.e. needs 3 evaluations, while needs 4). This is handled by the HyperPlonk SumCheck protocol with a fixed interpolation step before the MLE tables are reduced. While these computations are expensive, many of these operations can be performed in parallel. All polynomials can compute their evaluations independently, and within a polynomial, evaluations at different binary values can also be parallelized. Thus, HyperPlonk’s SumChecks have high compute and degrees of parallelism.

### v2

@@ Abstract.
Zero-Knowledge Proofs (ZKPs) are a rapidly growing technique for privacy-preserving and verifiable computation. ZKPs enable one party (a prover: ) to prove to another (a verifier: ) that a statement is true or correct without revealing any additional information. This powerful capability has led to ZKPs being applied and proposed for application in blockchain technologies, verifiable machine learning, and electronic voting. However, ZKPs have yet to see widespread, ubiquitous adoption due to the exceptionally high computational complexity of the proving process. Naturally, there has been recent work to accelerate ZKP primitives and protocols using GPUs and ASICs. However, the protocols considered so far face one of two challenges: they require a trusted setup for each new application or generate large proofs with high verification costs, limiting their applicability in scenarios with numerous verifiers or strict verification time constraints. HyperPlonk is a state-of-the-art ZKP protocol that supports both one-time, universal setup and small proof sizes/verification costs expected by publicly verifiable, consensus-based systems (e.g., blockchain). While HyperPlonk’s setup and verifier properties are highly desirable, the proving phase is costly. A HyperPlonk prover must compute on large bitwidths (e.g., 255-381b) and polynomials (e.g., of degree 224), employs computationally (e.g., MSM) and bandwidth (e.g., SumCheck) intensive kernels, and the complete protocol comprises many steps, each constituting distinct kernels. We present an accelerator, zkSpeed, to address these challenges and effectively accelerate HyperPlonk. zkSpeed provides hardware support for all major primitives (e.g., SumCheck and Multi-Scalar Multiplications (MSMs)) and judiciously schedules each protocol phase onto the allocated hardware. We leverage high-level synthesis to thoroughly explore and optimize the hardware design tradeoffs of each unit. These are then input into a full-chip simulator for large-scale design space exploration to optimize all aspects of the architecture in unison. Our Pareto analysis demonstrates that with a 366mm2 chip and 2 TB/s of off-chip bandwidth, zkSpeed is able to accelerate the entire proof generation by 801 (geomean) over a CPU baseline.
@@ 1. Introduction
Zero-Knowledge Proofs (ZKPs) are a privacy-preserving computation technique that enable one party (a prover ) to prove to another (a verifier ) that a computation (possibly with secret inputs) was performed correctly. The proof itself reveals no information about the (secret) inputs nor the function, and it can be verified much faster than the compute time of the original computation. Proof systems can be characterized by several important criteria, including: prover time, verifier time, proof size, protocol setup, and cryptographic assumptions required for the proof to be secure. Each deployment of ZKPs can result in a different set of criteria requirements. For example, private transactions (Ben-Sasson et al., 2014) require one proof per transaction be posted on a blockchain and distributed to all blockchain nodes, thus prioritizing small proof size.
@@
Recent years have seen significant advances and heterogeneity in proof systems (Groth, 2016; Gabizon et al., 2019; Chen et al., 2022; Xie et al., 2022; Setty, 2020; Abbaszadeh et al., 2024). Each optimizes for a distinct set of criteria constraints, and we note that in all systems, prover time is vital to optimize for, as faster provers enable new applications and proving is much slower than verification (lib, 2018; Chen et al., 2022). To understand the ZKP space, consider Groth16 (Groth, 2016) and Orion (Xie et al., 2022), two ZKPs with distinct characteristics that have also recently been accelerated (Zhang et al., 2021; Daftardar et al., 2024; Samardzic et al., 2024). Groth16 produces very short proofs (192 bytes) and enables millisecond verification latency, independent of proved computation size. However, it relies on a strong cryptographic assumption: a circuit-specific trusted setup where a trusted party generates keys using secret randomness. If this randomness is not properly discarded or the party is malicious, the system’s security is compromised, allowing false statements to be proven. Hence, ZCash (Ben-Sasson et al., 2014) and other applications have moved away from circuit-specific trusted-setup protocols (Nelson, 2022). On the other hand, Orion does not require a trusted setup or rely on computationally expensive elliptic curve cryptography, enabling a faster prover—but at the cost of large (8 MB) proofs. To put this in perspective, Orion’s proofs are roughly 4 larger than the maximum block size observed in Ethereum cryptocurrency (Monnot, 2024). Thus, a private transaction with an Orion proof would not fit into a single Ethereum block.
HyperPlonk (Chen et al., 2022) is a recent system gaining attention for the ZKP criteria it optimizes for, which are desirable in many applications. It has small proofs (typically 5 KB), low verification complexity, and supports universal trusted setup (Groth et al., 2018), which runs once for all time and then can be reused by new applications. A major contribution of HyperPlonk is its elimination of the Number Theoretic Transform (NTT), which is used by Groth16 and others, and instead its usage of SumCheck, which lowers the asymptotic runtime from (NTT) to (SumCheck). Here, represents the size of the underlying computation, and thus a lower runtime implies larger problems may be applied to ZKPs; something we cannot do today due to long runtime. Despite the improved runtime, the proving phase of HyperPlonk is still slow, running on the order of minutes to hours for some applications.
Accelerating HyperPlonk presents many challenges. First, the prover works over large, computationally demanding data structures, with polynomials on the order of 217-224 and bitwidths ranging from 255-381. Second, the protocol constitutes four major phases of computation, each comprising a different set of kernel functions. Kernels range from small and large compute-bound problems (e.g., SHA3 and MSM, respectively) to memory-bound (e.g., SumCheck). Compared to more traditionally accelerated kernels (e.g., NTT, matrix multiply, and convolution), HyperPlonk’s kernels are also less well understood. Finally, there is heterogeneity in data access patterns across kernels and phases. Some data structures are reused, either between or within kernels, and some are not. This raises questions on how to connect and schedule kernels as well as how to stage data to overlap communication with computation. The complexity of HyperPlonk requires careful design, new hardware optimizations, and thorough design space exploration to identify architectures with area-performance tradeoffs that both meet the computational needs of applications and justify the costs of hardware acceleration.
In this paper we present zkSpeed, a modular HyperPlonk accelerator that overcomes the challenges outlined above. zkSpeed comprises eight unique kernel accelerators, called units. Each unit supports required operators (e.g., large-word modular arithmetic) natively in hardware over a programmable problem size. The units collectively implement the complete zkSpeed chip architecture, which includes shared and unit-local scratchpads to capture reuse, capable of processing entire HyperPlonk proofs. Units and scratchpads communicate via a multi-channel shared bus, and we propose a schedule of each (sequential) HyperPlonk protocol phase onto zkSpeed accordingly. The schedule balances off-chip bandwidth constraints with on-chip reuse from scratchpads, unit computational throughput, and bus resource contention with a goal of overlapping off-chip communication with computation whenever possible. All unit designs and bandwidth constraints are configurable and can be tailored to use the least amount of area needed to achieve this goal. To optimize all units under a given off-chip bandwidth, we developed a zkSpeed simulator capable of sweeping over a wide range of design parameter settings. We consider thousands of unique zkSpeed designs. Analyzing the resulting Pareto frontier, we conclude that a 366mm2 chip with 2 TB/s of off-chip bandwidth can speedup proof generation by 801 (geomean) over a CPU baseline.
This paper makes the following contributions:
@@
A high-throughput, fully-pipelined accelerator to handle three flavors of SumCheck for HyperPlonk-based proofs.
@@
A novel implementation for modular inversion to compute fraction polynomials not seen in other ZKP protocols.
@@
Additional optimizations–including resource sharing of modular multipliers across units and on-chip memory compression schemes–that save upwards of 50% area per unit and up to 85% in bandwidth utilization.
@@
A full-chip design that achieves 801 gmean speedup over CPU at iso-compute area.
@@ 2. Background
In this section, we describe state-of-the-art ZKPs and introduce the mathematical principles underlying the key kernels of HyperPlonk.
@@
Today, the state-of-the-art ZKP protocols are zero-knowledge Succinct Non-interactive Arguments of Knowledge (zkSNARKs). zkSNARKs have three properties: (i) zero-knowledge, i.e., the proof does not reveal any information about the secret witness ; (ii) succinct, i.e., the proof has a few hundreds of bytes; and (iii) non-interactive, i.e., sends the proof to in one exchange.
zkSNARKs like HyperPlonk use polynomials to encode the correct execution of the target program. Polynomials can encode complex computations and constraints in a compact form, reducing the computational burden on to a few checks, instead of requiring inspection of individual operations in the program. In HyperPlonk, the two most time-consuming kernels are SumCheck and Multi-Scalar Multiplication (MSM). The SumCheck kernel lets demonstrate knowledge of a polynomial by verifying that its properties hold over a large set of inputs without revealing, e.g., the polynomial’s coefficients. MSMs ensure that is bound to that polynomial, preventing manipulation of the polynomial in the proof.
@@ 2.2. The SumCheck Kernel
SumCheck (Thaler, 2022) is an interactive protocol between a prover and a verifier . demonstrates to that it has correctly computed the sum of a polynomial over the boolean hypercube, i.e., over all Boolean (0/1) assignments of its variables.
@@
then sends coefficients or evaluations of to . checks that , and if so, generates a random challenge and sends it to .
In Round 2, asks to prove that
which is simply an instance of SumCheck, except over a ()-variate polynomial. Thus, in Round 2, computes and returns
@@ 2.3. Multilinear Polynomials
is multilinear because the maximum degree of individual variables is one. ZKP protocols use a general representation of multilinear polynomials, shown below for a 2-variable multilinear polynomial:
This representation defines polynomials using evaluations at specific input points (e.g., ) rather than using coefficients of symbolic terms (e.g., ). Just as a degree- polynomial is uniquely determined by coefficients, it can also be uniquely defined by its values at distinct points. This representation then maps cleanly to hardware in the form of lookup tables. For example, for a 2-variable multilinear polynomial, the binary values of index into the table to get the corresponding evaluation. In general, a multilinear polynomial with variables can be stored in a lookup table of entries. As we will discuss in Section 3, HyperPlonk uses multilinear polynomials as building blocks for higher-degree polynomials on which we then perform SumCheck and other computations. In the rest of the paper, we will use the term MLE table to refer to these lookup data structures (MLE stands for “multilinear extensions” (Thaler, 2022); the word “extension” indicates that these polynomials can be evaluated at non-boolean, or extended, values).
@@ 2.4. The MSM Kernel
MSMs are dot products between a vector of scalars and a vector of 2D or 3D points on an elliptical curve, e.g., . MSMs are used in ZKPs to perform commitments. A commitment is a cryptographic primitive that binds a prover to a value without revealing it. In zkSNARK protocols like HyperPlonk (Chen et al., 2022), scalars are the polynomial evaluations stored in MLE tables. Computing a dot-product reduces polynomials to a single value, i.e., the commitment.
MSMs are a bottleneck in ZKP provers and recent work has focused on accelerating them on ASIC and GPU (Daftardar et al., 2024; Zhang et al., 2021; Ma et al., 2023; Liu et al., 2024a; Ji et al., 2024; Lu et al., 2022; Zhou et al., 2024; Liu et al., 2024c, b; Hirner et al., 2025). The bottleneck is due to the extremely expensive elliptic curve point multiplications that they use. To reduce this cost, MSMs use Pippenger’s algorithm (Pippenger, 1976), which performs point multiplications via several point additions (PADDs). PADDs are still expensive, typically tens of regular modular multiplications. In the context of ASIC accelerators, the state-of-the-art (SZKP (Daftardar et al., 2024)) presents a framework for building scalable MSM architectures.
@@ 3. The HyperPlonk Protocol
In this section, we describe how HyperPlonk encodes the computation being proven, how this encoding shapes the structure of the underlying polynomials, and the key steps in proof generation.
@@ 3.1. Plonk-based encodings
In ZKPs, the program being proven must be converted into a specific form before generating the proof. While most prior works use Rank-1 Constraint System (R1CS) (lib, 2018; Groth, 2016)—a series of sparse matrix-vector encodings—HyperPlonk adopts a Plonk-based polynomial structure (Gabizon et al., 2019). R1CS and Plonk encodings map all program computations into addition, multiplication, and boolean operations, with nonlinear functions (e.g., branches) resolved via bit-wise decompositions.
Plonk-based encodings map each operation in a program’s execution to a gate – similar in principle to gates seen in other privacy-preserving protocols (Mo et al., 2023, 2025b; Chillotti et al., 2020; Feng et al., 2024) – that supports addition, multiplication, conditionality, and equality checks, as shown in Equation 1:
This can be viewed as a ZKP-specific instruction set where each gate acts as an instruction performing basic arithmetic or logical operations. In Equation 1, terms are binary control signals for left, right, multiply, and output ports. The term represents a constant input. Terms represent the gate’s inputs, and is its output. These last three terms are collectively called witnesses. An addition operation, for instance, is implemented by setting , , and other control and constant inputs to . This yields: . That is, if and only if the addition is correctly performed.
Each operation in a program is mapped to a Plonk gate. A program with operations has Plonk gates connected to form a circuit. A -variate polynomial represents the entire circuit, with binary assignments to variables representing individual gates. For example, if , the circuit has gates; represents the first gate.
is constructed using multilinear polynomials representing every term from Eq. 1. For example, is the left control input polynomial, is the first data input polynomial, and so on. Each of these polynomials also takes as an index. For example, in a program that maps to 16 gates, will tell us whether or not the multiplication control signal is enabled in gate 2. tells us what the output data signal is at gate 15. All individual polynomials are encoded in this fashion, and can be stored as MLE tables. These MLE tables are populated from the program trace in software before running the HyperPlonk prover. The MLEs are then used by the steps of the HyperPlonk protocol, which we discuss in the following sections.
@@ 3.2. SumCheck on Plonk-based Encodings
The key difference between the SumCheck used in HyperPlonk and the example in Section 2.2 lies in the structure of the underlying polynomials. While the earlier example operates on a single polynomial of arbitrary degree, HyperPlonk applies SumCheck to a version of the Plonk polynomial in Equation 1, as well as other polynomials discussed in Section 4.1. These polynomials consist of the products of multilinear polynomials, with terms of varying degrees. Figure 1 shows an example of how a SumCheck of the product of three multilinear polynomials is computed.
In the example from Section 2.2, if was multilinear, we would only need the evaluations at and because the result of summing over all variables yielded a univariate degree-1 polynomial. In this example, however, after summing over all variables, we are left with a univariate degree-3 polynomial. As mentioned in Section 2.3 (and from basic polynomial principles), a degree-3 polynomial must be evaluated at 4 unique points to fully characterize the polynomial. In the figure example, in the blue region, we are iterating over to compute the evaluations. For each iteration (red region), we must therefore evaluate each polynomial at . (Here, are the extensions that we compute from the existing evaluations at .) Then, in the green region, we compute the product for across all polynomials, and then sum across iterations. This is repeated for . After the summations, each polynomial’s MLE table is updated with a random challenge from the verifier. We refer to this step as MLE Update. For every boolean hypercube instance (e.g., in Figure 1), we extrapolate the linear polynomial in to the value of . For example, a table (for round 2) can be constructed from a table (from round 1) using the formula
The sum of all of these core computational steps is more expensive than baseline SumCheck. An additional layer of complexity is that HyperPlonk’s SumCheck polynomials exhibit degree variation across terms. This leads to imbalance in how many total evaluations are needed per term (i.e., needs 3 evaluations, while needs 4). This is handled by the HyperPlonk SumCheck protocol with a fixed interpolation step before the MLE tables are updated. While these computations are expensive, many of these operations can be performed in parallel. All polynomials can compute their extensions independently, and within a polynomial, extensions at different values can be parallelized. Thus, HyperPlonk’s SumChecks have high compute and degrees of parallelism.
SumCheck example illustrating the computation for an example polynomial.
HyperPlonk’s protocol steps, zKSpeed’s architecture, and dataflow for each protocol step.